OpenStudy (anonymous):

find three consecutive positive even integers such that the product of the second and third integers is twenty more than ten times the first integer

OpenStudy (mertsj):

The integers are: x, x+2, x+4 \[(x+2)(x+4)=10x+20\] \[x^2+6x+8=10x+20\] \[x^2-4x-12=0\] \[(x-6)(x+2)=0\] \[x=6, x=-2\] The integers are; 6,8,10 or -2,0,2

OpenStudy (anonymous):

amazing. thnx:)